# Diffuson-Dominated Thermal Transport Crossover from Ordered to Liquid-like Cu3BiS3:The Negligible Role of Ion Hopping

### v1

@@ 3 Conclusion
J.Y. acknowledges the National Natural Science Foundation of China through grants No. 52072188 and No. 12204254, the Program for Science and Technology Innovation Team in Zhejiang through grant No. 2021R01004, as well as the Institute of High-pressure Physics of Ningbo University for its computational resources. X.S. acknowledges funding supported by National Youth Talent Project and the Fundamental Research Funds for the Central Universities (D5000250021), and the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101034329 and the WINNING Normandy Program supported by the Normandy Region. S.L. acknowledges the computation resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at Sigma and Tetralith Clusters partially funded by the Swedish Research Council through grant agreements No. 2022-06725 and No. 2018-05973, as well as by the Vienna Scientific Cluster (VSC) in Austria.

### v2

@@ Acknowledgements
J.Y. acknowledges the National Natural Science Foundation of China through grants No. 52072188 and No. 12204254, the Program for Science and Technology Innovation Team in Zhejiang through grant No. 2021R01004, as well as the Institute of High-pressure Physics of Ningbo University for its computational resources. X.S. acknowledges funding supported by National Youth Talent Project and the Fundamental Research Funds for the Central Universities (D5000250021), and the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101034329 and the WINNING Normandy Program supported by the Normandy Region. X.S, K.M, and E.G gratefully thank Jean-François Lohier, Fabien Veillon, Christelle Bilot, and Jerôme Lecourt for technical support. S.L. acknowledges the computation resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) – at supercomputer centers NSC and PDC, partially funded by the Swedish Research Council through grant agreement no. 2022-06725 – and by the Vienna Scientific Cluster (VSC) in Austria.
Conflict of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.